what is the slope intercept form of the line described?
through: (3,5) parallel to y=7/4x+1

Answer by mananth(16949)

1 y = 7/4 x + 1
Divide by 1
y = 7/4 x + 1
Compare this equation with y=mx+b
slope m = 7/4
The slope of a line parallel to the above line will be the same
The slope of the required line will be 1 3/4
m= 7/4 ,point ( 3 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = 21/ 4 + b
b= -1/4
m= 7/4
Plug value of the slope and b in y = mx +b
The required equation is y = 7/4x- 1/ 4
m.ananth@hotmail.ca